OpenStudy (anonymous):

What is the simplified form of sqrt 162x^5y^6?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{162x^5y^6}\] \[\Large \sqrt{81*2*x^4*x*y^6}\] \[\Large \sqrt{81x^4y^6*2x}\] \[\Large \sqrt{(9x^2y^3)^2*2x}\] \[\Large \sqrt{(9x^2y^3)^2*2x}\] \[\Large \sqrt{(9x^2y^3)^2}*\sqrt{2x}\] \[\Large 9x^2y^3\sqrt{2x}\] So \[\Large \sqrt{162x^5y^6}\] simplifies to \[\Large 9x^2y^3\sqrt{2x}\] In other words \[\Large \sqrt{162x^5y^6} = 9x^2y^3\sqrt{2x}\]

jimthompson5910 (jim_thompson5910):

Where each variable is not negative.

OpenStudy (anonymous):

thanks sooooo much!

jimthompson5910 (jim_thompson5910):

you're welcome